OpenStudy (anonymous):

When surveyed, 5 out of every 8 students liked the food in the cafeteria. If 60 liked the food, how many were surveyed?

OpenStudy (slaaibak):

60 / 5 * 8 = answer

OpenStudy (anonymous):

do u have multi choise

OpenStudy (anonymous):

no I think it 96 but not sure

OpenStudy (anonymous):

ok hold up

OpenStudy (anonymous):

its 5/8 = 60/x

OpenStudy (anonymous):

there u go

OpenStudy (slaaibak):

96 is correct.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

could that be

OpenStudy (slaaibak):

96 * 5/8 = 60

OpenStudy (anonymous):

ur right